A ball thrown vertically upward from ground level hits the ground after 4seconds calculate the maximum height it reached during its journey
1
Expert's answer
2019-10-14T09:43:27-0400

The time of motion to the highest point is 2 s. So, the initial velocity of a ball


"v_i=gt=9.8\;\rm m/s\times 2\; s=19.6\; m/s"

The maximum height that ball reached


"h_{max}=\frac{v_i^2}{2g}=\frac{(19.6\:\rm m/s)^2}{2\times 9.8\:\rm m/s^2}=19.6\:\rm m"